OpenStudy (anonymous):

negative square root of 48b^4

OpenStudy (jdoe0001):

well, just get the \(\sqrt{48b^4}\) and slap a "-" in front of it :)

OpenStudy (anonymous):

do u know the answer to it please reply if u do or nnot thanks -.-

OpenStudy (anonymous):

how to do the square root sign on keyboard?

OpenStudy (anonymous):

\[negative squareroot of 48b^4=-\sqrt{48b ^{4}}=\sqrt{16*3b ^{4}}=-4\sqrt{3}b ^{2}\]

OpenStudy (anonymous):

omg thank you so much i think i got the right answer :)

OpenStudy (anonymous):

click equation and then click 4th row and then click square root x

OpenStudy (anonymous):

\[^{5}\]

OpenStudy (anonymous):

\[^{7}\]

OpenStudy (anonymous):

itwont happenn